# Diffusion Transformer Policy

### v1

@@ 1 Introduction
Traditional robot learning paradigm usually relies on large-scale data collected for a specific robot and task, but collecting robot data for generalist tasks is time-consuming and expensive due to the limitations of robot hardware in the real world. Nowadays, the foundational models OpenAI (2022; 2023; 2021); Rombach et al. (2021) in Natural Language Process and Computer Vision, pretrained on broad, diverse, task-agnostic datasets, have demonstrated powerful ability in solving downstream tasks either zero-shot or with a few task-specific samples. It is principally possible that a general robot policy exposed to large scale diverse robot datasets improves generalization and performance on downstream tasks Brohan et al. (2022; 2023). However, it is challenging to train a general robot policy on a large scale of cross-embodiment datasets with diverse sensors, action spaces, tasks, camera views, and environments.
@@
Toward a unified robot policy, existing works directly map visual observation and language instructions to actions with large visual-language-action models for robot navigation Shah et al. (2023a; b) or manipulation Brohan et al. (2022; 2023); Kim et al. (2024); Team et al. (2024), and demonstrate zero-shot or few-shot generation to new environments. Robot Transformers Brohan et al. (2022; 2023); Padalkar et al. (2023) present robot policy based on transformer architecture, and demonstrate robust generalization by training on the large scale of Open X-Embodiment Dataset Padalkar et al. (2023). Octo Team et al. (2024) follows the autoregressive transformer architecture with a diffusion action head, while OpenVLA Kim et al. (2024) discretizes the action space and leverage the pretrained visual-language model to build VLA model exposed to Open X-Embodiment Dataset Padalkar et al. (2023). Though those Visual-Language-Action (VLA) models Team et al. (2024); Kim et al. (2024) have shown the potential to learn robot policy from the large cross embodiment datasets Padalkar et al. (2023), the diversity of robot space among the cross embodiment datasets still limits the generalization of previous models with autoregressive transformers.
Recent diffusion policy Chi et al. (2023); Ze et al. (2024); Ke et al. (2024) has shown its stable ability in robot policy learning for single task imitation learning with UNet or cross attention architecture, and diffusion transformer demonstrates its scalability in multi-modal image generation Peebles & Xie (2023). Specifically, Octo Team et al. (2024) presents a generalist policy that denoises the action with a small MLP network conditioned on a single embedding of auto-regressive multi-modal transformer. However, the robot space of large-scale cross-embodiment datasets contains various cameras views and diverse action spaces, which poses a significant challenge for a small MLP to denoise the continuous action conditioned on a single action head embedding.
In this paper, we design a Diffusion Transformer architecture for generalist robot policy learning. Similar to previous robot transformer models Brohan et al. (2022; 2023); Padalkar et al. (2023); Team et al. (2024); Kim et al. (2024), we leverage the transformer as our base module to retain the scalability on the large-scale cross-embodiment datasets. Different from Brohan et al. (2022; 2023); Padalkar et al. (2023); Team et al. (2024); Kim et al. (2024), we present an in-context conditional diffusion transformer architecture to denoise the action chunks, rather than utilizing a small MLP to denoise action embedding to continuous actions as illustrated in Figure 1. The Diffusion Transformer Policy retains the scalability of transformer for diffusion, and thus more effectively generalizes action policy from the large diverse datasets.
In a nutshell, we present a Diffusion Transformer Policy, that incorporates a causal transformer as an in-context conditional diffusion backbone and denoise continuous action chunks with the transformers. Extensive experiments demonstrate Diffusion Transformer Policy achieves considerably better performance on two large-scale Sim datasets, Maniskill2 and Calvin, compared to corresponding diffusion action head Team et al. (2024) and discretized action head Brohan et al. (2022) baselines. Meanwhile, The proposed model trained on the Open X-Embodiment Dataset achieves better generalization performance compared to the baseline methods on the Real Franka platform.
@@
Figure 1: Illustrations of different robot policy architectures. (a) is the common robot transformer architecture with discretization actions, e.g., Robot Transformer Brohan et al. (2022; 2023) and OpenVLA Kim et al. (2024). (b) is the transformer architecture with diffusion action head which denoises the continuous action with a small MLP with embedding from the causal transformer, e.g., Octo Team et al. (2024). (c) is the proposed Diffusion Transformer architecture that utilizes the large transformer to denoise actions in an in-context conditioning style.
@@ 2 Related Work
Diffusion Policy Denoise diffusion techniques Ho et al. (2020); Rombach et al. (2022) are pioneering image generation, and recent Diffusion Policy Chi et al. (2023); Ze et al. (2024); Ke et al. (2024) has exhibited a powerful ability in modeling multimodal actions compared to previous robot policy strategies in both 2D Chi et al. (2023) and 3D observations Ze et al. (2024); Ke et al. (2024). Current diffusion policy approaches usually follow an Unet structure or a shallow cross-attention network for a single manipulation task, leaving large-scale multimodal diffusion policy poorly investigated. For example, 3D diffusion Policy Ze et al. (2024) presents a diffusion approach conditional on a 3D point cloud, while 3D diffuser actor Ke et al. (2024) proposes a 3D diffusion strategy based on point cloud with cross-attention. Differently, we present a scalable in-context conditioning diffusion transformer architecture for generalist robot policy. Recent generalist policy Octo Team et al. (2024) conditions the denoise process on the embedding from the Transformer model with a small MLP diffuser. By contrast, the diffuser in Diffusion Transformer Policy is a large Transformer architecture.
Generalist Robot Policies The embodiment community has shown increasing interest in generalist robot policy with foundational multi-modal models for both robot navigation Shah et al. (2023a; b); Yang et al. (2024); Sridhar et al. (2024); Huang et al. (2023) and manipulation Bousmalis et al. (2023); Brohan et al. (2022); Shah et al. (2023c); Shridhar et al. (2023); Brohan et al. (2023); Padalkar et al. (2023); Kim et al. (2024); Team et al. (2024). Recent approaches Brohan et al. (2022; 2023); Padalkar et al. (2023); Kim et al. (2024); Team et al. (2024) aim to achieve generalist policy with scalable Visual-Language-Action models. We follow this paradigm to approach generalist and adaptive robot policy.  Brohan et al. (2022; 2023); Padalkar et al. (2023); Kim et al. (2024) construct the action token by discretizing each dimension of the robot actions separately into 256 bins. However, this discretization strategy incurs internal deviation in robot execution. Unlike those methods, we present a Diffusion Transformer Generalist Policy, which denoises the continuous actions with a large Transformer model. The proposed approach retains the scalability of the Transformer and meanwhile facilitates the modeling of cross-embodiment action chunk representations. Meanwhile, the Diffusion Transformer Policy aligns robot action together with the language instructions and image observations as an in-context conditional style.
@@ 3 Method
We describe the proposed architecture of diffusion transformer policy in this section, a DiT-based generalist diffusion policy model that can be adapted to new environments and embodiment. The model, built from a diffusion transformer architecture, achieves better generalization on both novel camera views and environments from large amounts of diverse robot data.
@@ 3.1 Architecture
Instruction Tokenization. The language instructions are tokenized by a frozon CLIP Radford et al. (2021) model.
Image observation Representations. The image observations first pass into the DINOv2 Oquab et al. (2023) to obtain the image patch features. Note that DINOv2 Oquab et al. (2023) is trained on the web data which is different from the robot data, we thus jointly optimize the DINOv2 parameters together with Transformers through an end-to-end way.
Q-Former. To reduce the computation cost, a Q-Former together with FiLM Perez et al. (2018) conditioning is incorporated to select image features from the patch features of DIDOv2 Oquab et al. (2023) by instruction context.
Action Tokenization. We use the end-effector action and represent each action with a 7D vector, including 3 dimensions for translation vector, 3 dimensions for rotation vector, and a dimension for gripper position. To align the dimension with image and language tokens, we simply pad the action vector with zeros to construct the action token. We only add the noise into the 7D action vector during denoise diffusion optimization.
Architecture. Our core design is the Diffusion Transformer structure Peebles & Xie (2023) which denoises action token chunks, instead of each single action token, conditioned on image observation and instruction tokens by an in-context conditioning style with a causal transformer network, i.e., we simply concatenate image tokens, language tokens, and timestep token in the front of the sequence, equally treating the noisy action tokens from the image/instruction tokens as illustrated in Figure 2. This design retains the scaling properties of transformer networks. The model, conditioned on language instructions and image observations with the causal transformer structure, is supervised by the noise that we add to the continuous actions. In other words, we conduct the diffusion objective directly in the action chunk space with a large transformer model, differently from a diffusion action head with a few MLP layers Team et al. (2024).
The proposed Diffusion Transformer Policy is a general design that can be scaled to different datasets, and demonstrate excellent performance. Meanwhile, we can also add additional observation tokens and input into the transformer structure. Appendix A provides more details.
@@ 3.2 Training Objective
In our architecture, the denoising network $\epsilon_{\theta}({\bm{x}}^{t},c_{obs},c_{instruction},t)$ is the whole causal transformer, where $c_{obs}$ is the image observation, $c_{instruction}$ is the language instruction, and $t\in{1,2,...T}$ is step index in our experiments. During the training stage, we sample a Gaussian noise vector ${\bm{x}}^{t}\in\mathcal{N}(\mathbf{0},{\bm{I}})$ at timestep $t$, where $T$ is the number of denoising timesteps, and add it to action ${\bm{a}}$ as $\hat{{\bm{a}}}$ to construct the noised action token, finally predicting the noise vector $\hat{{\bm{x}}}$ based on the denoising network $\epsilon_{\theta}(\hat{{\bm{a}}},c_{obs},c_{instruction},t)$, where $t$ is randomly sampled during training. We optimize the network with MSE loss between ${\bm{x}}^{t}$ and $\hat{{\bm{x}}^{t}}$.
@@
${\bm{x}}^{t-1}=\alpha({\bm{x}}^{t}-\gamma\epsilon_{\theta}({\bm{x}}^{t},c_{obs},c_{instruction},t)+\mathcal{N}(\mathbf{0},\sigma^{2}{\bm{I}})).$
where $\alpha$, $\gamma$, $\sigma$ is the noise scheduler Ho et al. (2020). In our experiments, $\epsilon_{\theta}$ is to predict the noise that adds to the action.
@@ 3.3 PreTraining Data
To evaluate the proposed Diffusion Transformer Policy, we choose Open X- Embodiment datasets Padalkar et al. (2023) for pretraining the model. We mainly follow Team et al. (2024); Kim et al. (2024) to choose the datasets and set the weights for each dataset. We normalize the actions similarity to Padalkar et al. (2023) and filter out outlier actions in the dataset. Additional details are provided in Appendix B.
@@ 3.4 Pretraining Details
We devise the proposed Diffusion Transformer architecture and evaluate the pretraining approach in the large cross-embodiment datasets Padalkar et al. (2023). We use the DDPM Ho et al. (2020) diffusion objective in the pretraining stage with $T=1000$ for the Open X-Embodiment dataset Padalkar et al. (2023), while we set $T=100$ with DDIM Song et al. (2020) for zero-shot evaluation to accelerate the inference. According to the preliminary experiment from Maniskill2 Gu et al. (2023), we use 2 observation images and predict 32 action chunks. We filter out the action chunks that include outlier values. We train the network with AdamW Loshchilov (2017) by 100,000 steps. We set the learning rate of the casual transformer and Q-Former as 0.0001, the learning rate of the pretrained image tokenizers as 0.00001, and the batch size as 8902. More pretraining details are provided in the Appendix.
Figure 2: Our model is a Transformer diffusion structure. The model first incorporates a pretrained CLIP network to obtain instruction tokens. Meanwhile, we use the DINO-V2 Oquab et al. (2023) model to encode image observations, followed by a Q-Former to query observation features for each image observation. Next, we concatenate instruction tokens, image observation tokens, timestep token, and noised action tokens together to construct a token sequence as the input for transformer network to denoise the raw actions.
@@ 4 Experiments
We evaluate the proposed methods with two baselines in three environments. We leverage Maniskill2 to present the ability of Diffusion Transformer Policy on large scale novel view generalization. Meanwhile, we demonstrate the generalization of the pretrained Diffusion Transformer Policy on CALVIN benchmark. Lastly, we further show the generalization of Diffusion Transformer Policy on Real Franka Arm.
@@ 4.1 Baselines
Discretization Action Head We implement the RT-1 Brohan et al. (2022) style baseline models with a similar structure as ours. We maintain the Instruction Tokenization and Image Tokenization strategy in the proposed method. Different from ours, we follow RT-1 Brohan et al. (2022) to discretize each dimension of the action into 256 bins, and leverage the transformer network to predict the action bin indexes. Following Brohan et al. (2022; 2023), we use cross-entropy loss to optimize the network.
Diffusion Action Head. We also implement a diffusion action head strategy  Team et al. (2024) in our architecture. Specifically, we utilize a three-layer MLP network as our denoising network condition on the output of each action token embedding by the same transformer architecture as ours.
4.2 Maniskill2
Maniskill2 Gu et al. (2023) is the next generation of the SAPIEN Maniskill benchmark Mu et al. (2021), which is widely used to evaluate the generalized manipulation ability of the embodied models. It contains 20 different manipulation tasks families and over 4 million demonstration frames with different settings, including rigid/soft body, single/dual arm, etc. Maniskill2 also provides a fast and easy way to change the camera view and replay the trajectories. It is useful for the researchers working on generalized policy.
Setup. In our experiments, we select 5 tasks (PickCube-v0, StackCube-v0, PickSingleYCB-v0, PickClutterYCB-v0, PickSingleEGAD-v0) from Maniskill2, and then construct a camera pool with 300,000 random cameras, then sample 20 cameras from the camera pool to render a trajectory each time, and finally obtain about 40K trajectories totally. Moreover, we split the dataset into training set and validation set according to a ratio of 19:1. During the splitting, it is guaranteed that the single trajectory rendered under different camera views will appear in either training set or validation set in order to avoid data leaking. Specially, there are 74 different categories to pick and place in the task family PickSingleYCB-v0. In addition to the mentioned rules, we ensure each category can be found in both training set and validation set. After that, we sample 100 trajectories for each task family randomly from the validation set, constructing a close loop evaluation dataset with 500 trajectories in total. While training, considering the balance between different task families, we adjust the number of data pieces to the same by simply copying the trajectories from task family with fewer trajectories originally. The ability of the model is measured by the success rate of each task family while executing the close loop evaluation dataset.
Optimization Details. We optimize the network with AdamW Loshchilov (2017) by 50,000 steps on Maniskill2 and we set the learning rate as 0.0001. The number of training timesteps $T$ is 100 in Maniskill2 and the global batch size is 1024.
Comparisons. Table 1 compares the proposed method with discretized action head Brohan et al. (2022; 2023) and diffusion action head Team et al. (2024). Here, the proposed method use the same backbone and transformer architecture as the baseline methods. The experiments demonstrate Diffusion Transformer Policy achieves better results compared to Discretization Action Head strategy Brohan et al. (2022) under the large scale novel view maniskill2 benchmark. Specifically, we observe the proposed Diffusion Transformer Policy achieves clear better performance compared to the baselines. Meanwhile, Diffusion Transformer Policy demonstrates better performance in more complex tasks, e.g., Diffusion Transformer Policy improves diffusion action head Team et al. (2024) by 20% in task PickSingleYCB and by 12% task PickClutterYCB. Those experiments show that Diffusion Transformer Policy achieves better scalability in the large scale diverse datasets, and meanwhile achieves better generation in camera view generalization.
Table 1: Comparision with two baseline methods on Maniskill2 under success rate. SingleYCB indicates PickSingleYCB, ClutterYCB indicates PickClutterYCB, SingleEGAD indicates PickSingleEGAD. Disc ActionHead indicates Discretized Action Head strategy Brohan et al. (2022), while Diff ActionHead shows Diffusion Action Head Team et al. (2024).
Method All PickCube StackCube SingleYCB ClutterYCB SingleEGAD
Disc ActionHead 30.2% 41.0% 33.0% 22.1.0% 54.0%
Diff ActionHead 58.6% 86.0% 76.0% 37.0% 24.0% 70.0%
DiT Policy(ours) 65.8% 79.0% 80.0% 62.0% 36.0% 72.0%

### v6

@@ 1 Introduction
Figure 1: Comparisons between previous state-of-the-art approaches and ours on popular generalist vision-language-action simulation benchmarks. LIBERO [36] and Calvin [43] are for finetuning generalization, while SimplerEnv [34] is to evaluate the generalization to environment variances under Real-to-Sim Benchmark.
@@
Traditional robot learning paradigm usually relies on large-scale data collected for a specific robot and task, but collecting robot data for generalist tasks is time-consuming and expensive due to the limitations of robot hardware in the real world. Nowadays, the foundational models [47, 48, 46, 58, 37] in Natural Language Process and Computer Vision, pretrained on broad, diverse, task-agnostic datasets, have demonstrated powerful ability in solving downstream tasks either zero-shot or with a few task-specific samples. It is principally possible that a general robot policy exposed to large scale diverse robot datasets improves generalization and performance on downstream tasks [6, 7]. However, it is challenging to train a general robot policy on a large scale of cross-embodiment datasets with diverse sensors, action spaces, tasks, camera views, and environments.
@@
Toward a unified robot policy, existing works directly map visual observation and language instructions to actions with large vision-language-action models for robot navigation [62, 63] or manipulation [6, 7, 28, 69], and demonstrate zero-shot or few-shot generalization to new environments. Robot Transformers [6, 7, 50] present robot policy based on transformer architecture, and demonstrate robust generalization by training on the large scale of Open X-Embodiment Dataset [50]. Octo [69] follows the autoregressive transformer architecture with a diffusion action head, while OpenVLA [28] discretizes the action space and leverage the pretrained vision-language model to build VLA model exposed to Open X-Embodiment Dataset [50]. Though those Vision-Language-Action (VLA) models [69, 28] have shown the potential to learn robot policy from the large cross embodiment datasets [50], the diversity of robot space among the cross embodiment datasets still limits the generalization.
@@
Figure 2: Illustrations of different robot policy architectures. (a) is the common robot transformer architecture with discretization actions, e.g., Robot Transformer [6, 7] and OpenVLA [28]. (b) is the transformer architecture with diffusion action head which denoises the individual continuous action with a small network condition on each embedding from the causal transformer, e.g., Octo [69]. (c) is the proposed Diffusion Transformer architecture that utilizes the large transformer to denoise actions in an in-context conditioning style.
Recent diffusion policy [12, 77, 26, 57] has shown its stable ability in robot policy learning for single task imitation learning with UNet or cross attention architecture, and diffusion transformer demonstrates its scalability in multi-modal image generation [52]. Specifically, Octo [69] presents a generalist policy that denoises the action with a small MLP network conditioned on a single embedding of auto-regressive multi-modal transformer. However, the robot space of large-scale cross-embodiment datasets contains various cameras views and diverse action spaces, which poses a significant challenge for a small MLP to separately denoise each continuous action conditioned on a single action head embedding. Meanwhile, previous diffusion policies [69, 57, 26] first fuse historical image observations and instruction into embeddings before the denoising process, which might limit action denoising learning since action anticipation usually relies directly on detailed historical observations rather than fused embedding.
In this paper, we design a Diffusion Transformer architecture for generalist robot policy learning. Similar to previous robot transformer models [6, 7, 50, 69, 28], we leverage the transformer as our base module to retain the scalability on the large-scale cross-embodiment datasets. Different from [6, 7, 50, 69, 28], we present an in-context conditional diffusion transformer architecture to denoise the action chunks (i.e., action sequence), rather than utilizing a small shared MLP to separately denoise each action embedding to continuous actions as illustrated in Figure 2. Meanwhile, the in-context conditional design enables the action denoising learning directly condition on each image observation patches, which supports the denoising model in perceiving the subtle nuances (e.g., action delta) in historical visual observations. Also, DiT Policy retains the scalability of transformer for diffusion, allowing for more effective generalization from large and diverse embodiment datasets.
In a nutshell, we present a Diffusion Transformer Policy, that incorporates a causal transformer as an in-context conditional diffusion backbone and denoise continuous action chunks. Extensive experiments demonstrate DiT Policy achieves considerably better performance on two large-scale Sim datasets, Maniskill2 (novel camera views) and Calvin, compared to baselines. Meanwhile, the proposed model trained on the Open X-Embodiment Dataset achieves better generalization performance compared to Octo and OpenVLA on the Real Franka and Libero.
@@ 2 Related Work
Diffusion Policy Denoise diffusion techniques [21, 59, 15, 52, 8] are pioneering image generation, and recent Diffusion Policy [35, 72, 9, 71, 10, 12, 77, 26, 57] has exhibited a powerful ability in modeling multimodal actions compared to previous robot policy strategies. Current diffusion policy approaches usually follow an Unet structure or a shallow cross-attention network for a single manipulation task, leaving large-scale multimodal diffusion policy poorly investigated. For example, 3D diffusion Policy [77] denoises policy conditioned on a 3D point cloud, while 3D diffuser actor [26] proposes a 3D diffusion strategy based on point cloud with cross-attention. Differently, we present a scalable in-context conditioning diffusion transformer architecture, which directly conditions on each historical observation. Recent generalist policy Octo [69] conditions the denoise process on the embedding from the Transformer model with a small MLP diffuser. By contrast, the diffuser in Diffusion Transformer Policy is a large Transformer architecture.
Generalist Robot Policies Language-conditioned policy [56, 39, 19, 45, 78, 11] is more suitable and general for real applications, and the embodiment community has shown increasing interest in generalist robot policy with foundational multi-modal models for both robot navigation [62, 63, 76, 67, 23, 4] and manipulation [5, 6, 64, 65, 7, 50, 28, 69, 16, 17, 51, 56, 40, 74, 2, 75, 42, 25, 60]. Recent approaches [6, 7, 50, 28, 69] aim to achieve generalist policy with scalable Vision-Language-Action models. We follow this paradigm to approach generalist and adaptive robot policy.  [6, 7, 50, 28] construct the action token by discretizing each dimension of the robot actions separately into 256 bins. However, this discretization strategy incurs internal deviation in robot execution. Unlike those methods, we present a Diffusion Transformer Generalist Policy, which denoises the continuous actions with a large Transformer model. The proposed approach retains the scalability of the Transformer and meanwhile facilitates the modeling of cross-embodiment action chunk representations. Meanwhile, the Diffusion Transformer Policy aligns robot action together with the language instructions and image observations as an in-context conditional style.
@@ 3 Method
We describe the architecture of diffusion transformer policy in this section, a DiT-based generalist diffusion policy that can be adapted to new environments and embodiments.
Figure 3: Our model is a Transformer diffusion structure. The model first incorporates a pretrained CLIP network to obtain instruction tokens. Meanwhile, we use the DINO-V2 [49] model to encode image observations, followed by a Q-Former to query observation features for each image observation. Next, we concatenate instruction tokens, image observation features, timestep, and noised action together to construct a token sequence as the input for transformer network to denoise the raw actions.
@@ 3.1 Architecture
Instruction Tokenization. The language instructions are tokenized by a frozon CLIP [55] model.
Image observation Representations. The image observations first pass into the DINOv2 [49] to obtain the image patch features. Note that DINOv2 is trained on the web data which is different from the robot data, we thus jointly optimize the DINOv2 parameters together with Transformers through an end-to-end way.
Q-Former. To reduce the computation cost, a Q-Former [30] together with FiLM [53] conditioning is incorporated to select image features from the patch features of DINOv2 [49] by instruction context.
Action Preprocess. We use the end-effector action and represent each action with a 7D vector, including 3 dimensions for the translation vector, 3 dimensions for the rotation vector, and a dimension for the gripper position. To align the dimension with image and language tokens, we simply pad the continuous action vector with zeros to construct the action representation. We only add the noise into the 7D action vector during denoise diffusion optimization.
Architecture. Our core design is the Diffusion Transformer structure [52] which denoises action token chunks, instead of each single action token, conditioned directly on image observation and instruction tokens by an in-context conditioning style with a causal transformer, i.e., we simply concatenate image features, language tokens, and timestep embedding in the front of the sequence, equally treating the noisy action from the instruction tokens as illustrated in Figure 3. This design retains the scaling properties of transformer networks, and allows the denoising learning conditioned directly on image patches, thus facilitating the model capture the detailed action changes in historical observations. The model, conditioned on language instructions and image observations with the causal transformer structure, is supervised by the noise that we add to the continuous actions. In other words, we conduct the diffusion objective directly in the action chunk space with a large transformer model, differently from a shared diffusion action head with a few MLP layers [69].
The proposed Diffusion Transformer Policy is a general design that can be scaled to different datasets, and demonstrates excellent performance. Meanwhile, we can also add additional observation tokens and input into the transformer structure. Appendix A provides more details.
@@ 3.2 Training Objective
In our architecture, the denoising network $\epsilon_{\theta}({\bm{x}}^{t},c_{obs},c_{instru},t)$ is the entire causal transformer, where $c_{obs}$ is the image observation, $c_{instru}$ is the language instruction, and $t\in{1,2,...T}$ is the step index in our experiments. During the training stage, we sample a Gaussian noise vector ${\bm{x}}^{t}\in\mathcal{N}(\mathbf{0},{\bm{I}})$ at timestep $t$, where $T$ is the number of denoising timesteps, and add it to action ${\bm{a}}$ as $\hat{{\bm{a}}}$ to construct the noised action token, finally predicting the noise vector $\hat{{\bm{x}}}$ based on the denoising network $\epsilon_{\theta}(\hat{{\bm{a}}},c_{obs},c_{instru},t)$, where $t$ is randomly sampled during training. We optimize the network with MSE loss between ${\bm{x}}^{t}$ and $\hat{{\bm{x}}^{t}}$.
@@
${\bm{x}}^{t-1}=\alpha({\bm{x}}^{t}-\gamma\epsilon_{\theta}({\bm{x}}^{t},c_{obs},c_{instru},t)+\mathcal{N}(\mathbf{0},\sigma^{2}{\bm{I}})).$ (1)
where $\alpha$, $\gamma$, $\sigma$ is the noise scheduler [21]. In our experiments, $\epsilon_{\theta}$ is to predict the noise that adds to the action.
@@ 3.3 Pretraining Data
To evaluate the proposed Diffusion Transformer Policy, we choose Open X- Embodiment datasets [50] for pretraining the model. We mainly follow [69, 28] to choose the datasets and set the weights for each dataset. We normalize the actions similar to [50] and filter out outlier actions in the dataset. Additional details are provided in Appendix B.
@@ 3.4 Pretraining Details
We devise the proposed Diffusion Transformer architecture and evaluate the pretraining approach in the large cross-embodiment datasets [50]. We use the DDPM [21] diffusion objective in the pretraining stage with $T=1000$ for the Open X-Embodiment dataset [50], while we set $T=100$ with DDIM [66] for zero-shot evaluation to accelerate the inference. According to the preliminary experiment from Maniskill2 [18], we use 2 observation images and predict 32 action chunks. We train the network with AdamW [38] by 100,000 steps. We set the learning rate of the casual transformer and Q-Former as 0.0001, the learning rate of DINOv2 as 0.00001, and the batch size as 8902. More pretraining details are provided in the Appendix A.
Table 1: Comparison with RT-1-X [6], Octo-base [69] and OpenVLA-7B [28] on SimplerEnv (average variance and matching results of Google Robot [6]). The results are reported as success rate.
Method coke__\__can move__\__near drawer
match variant match variant match variant
RT-1-X [6] 56.7% 49.0% 31.7% 32.3% 59.7% 29.4%
Octo-Base [69] 17.0% 0.6% 4.2% 3.1% 22.7% 1.1%
OpenVLA-7B [28] 16.3% 54.5% 46.2% 47.7% 35.6% 17.7%
DiT Policy(Ours) 72.7% 60.0% 56.7% 57.5% 46.3% 37.5%
@@ 4 Experiments
We evaluate the proposed methods with two baselines in three environments. We leverage Maniskill2 to present the ability of Diffusion Transformer Policy on large scale novel view generalization. Meanwhile, we demonstrate the generalization of the pretrained Diffusion Transformer Policy on CALVIN benchmark. Lastly, we further show the generalization of DiT Policy on Real Franka Arm.
@@ 4.1 Baselines
Discretization Action Head We implement the RT-1 [6] style baseline models with a similar structure as ours. We keep the Instruction Tokenization and Image backbone. Different from ours, we discretize each dimension of the action into 256 bins [6], and leverage the transformer network to predict the action bin indexes. Following [6, 7], we use cross-entropy loss to optimize the network.
Diffusion Action Head We also implement a diffusion action head strategy  [69]. Specifically, we utilize a three-layer MLP network as our denoising network condition on the output of each action token embedding by the same transformer architecture as ours. Notably, this baseline has more parameters (the additional MLP) compared to DiT policy.
4.2 SimplerEnv
SimplerEnv [34] is a Real-to-Sim platform for evaluating the policy learned from real robot data with a simulation platform. In this section, we compare with the popular generalist policy trained on Open X-embodiment dataset, including RT-1-X [6, 50], Octo [69] and OpenVLA [28], on Google Robot Simulation with different variances. We follow the evaluation from SimplerEnv [34] for a fair comparison, which includes “pick up coke can”, “move an object near to others”, “open drawer”, “close drawer”.
Table 1 demonstrates that the proposed approach achieves a strong generalization performance under different variances, including background, texture, objects, spatial positions, and so on. The variance experiments show the robust of Diffusion Transformer Policy.